find the coordinates of the vertices of triangle ABC formed by the intersection of the lines x+y=0, x=0, y=x-1. Hence find the area of triangle ABC?
Answer by Emsley(35)

x+y=0 meets x=0 at (0,0) y=x-1 meets x=0 at (0,-1)
Solving x+y=0 and y=x-1 simultaneously sub y=x-1 into x+y=0 we get x+x-1=0
2x=1 x=1/2 Sub back to get y=-1/2
So points are (0,0) (0,-1) and (1/2,-1/2)